# Towards Efficient PCSEL Design: A Data-Driven Approach for Design Insights

### v3

@@ Conclusion
In this work, we developed a photonic crystal simulation framework based on Coupled-Wave Theory (CWT) to analyze and optimize the performance of Photonic Crystal Surface-Emitting Lasers (PCSELs). By systematically modeling the interaction of optical waves within a structured photonic lattice, we derived key performance metrics, including surface-emitting efficiency ($SEE$) and the quality factor ($Q$). To further enhance the design process, we implemented a neural network model trained on a large dataset of simulated PCSEL structures. of the photonic crystal, enabling rapid optimization of device performance.
Additionally, we employed Shapley value analysis to interpret the contributions of different Fourier components, providing insights into the structural features that most significantly impact laser efficiency. Our results demonstrate that machine learning can serve as a powerful tool for guiding photonic crystal design, accelerating the discovery of high-performance PCSEL configurations. The methodology proposed in this study is flexible and can be extended to different epitaxial structures and laser designs, paving the way for future advancements in high-efficiency surface-emitting lasers.
@@ Acknowledgements
This research was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62174144, the Shenzhen Science and Technology Program under Grants No. JCYJ20220818102214030, No. KJZD20230923115114027, No. JSGG20220831093602004, No. KJZD20240903095602004, the Guangdong Key Laboratory of Optoelectronic Materials and Chips under Grant No. 2022KSYS014, the Shenzhen Key Laboratory Project under Grant No. ZDSYS201603311644527, the Longgang Key Laboratory Project under Grants No. ZSYS2017003 and No. LGKCZSYS2018000015, and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300701), Hefei National Laboratory, Hefei 230088, China.
The authors would thank Ms. Floria Chen and Ms. Chuanyi Yang for their mentorship and revision tips. The authors would also thank Mr. Qi Xin and Mr. Sixuan Mao for their suggestions on frameworks of artificial neural networks.
@@
H.H. conceived the experiment and conducted the majority of the manuscript writing. H.H. and Z.X. conducted the experiment and analyzed the results. Z.X. wrote the section on the principles of the neural network and prepared Figure 1. Z.Z. provided guidance and funding support. All authors reviewed the manuscript.

### v4

@@ Conclusion
In this work, we developed a PCSELs simulation framework based on CWT to analyze and optimize the performance of PCSELs. By systematically modeling the interaction of optical waves within a structured photonic lattice, we derived key performance metrics, including surface-emitting efficiency ($SEE$) and the quality factor ($Q$). To further enhance the design process, we implemented a neural network model trained on a large dataset of simulated PCSEL structures. of the PhC, enabling rapid optimization of device performance.
Additionally, we employed Shapley value analysis to interpret the contributions of different Fourier components, providing insights into the structural features that most significantly impact laser efficiency. Our results demonstrate that machine learning can serve as a powerful tool for guiding PhC design, accelerating the discovery of high-performance PCSEL configurations. The methodology proposed in this study is flexible and can be extended to different epitaxial structures and laser designs, paving the way for future advancements in high-efficiency surface-emitting lasers.
@@ Acknowledgements
This research was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62174144, the Shenzhen Science and Technology Program under Grants No. JCYJ20220818102214030, No. KJZD20230923115114027, No. JSGG20220831093602004, No. KJZD20240903095602004, the Guangdong Key Laboratory of Optoelectronic Materials and Chips under Grant No. 2022KSYS014, the Shenzhen Key Laboratory Project under Grant No. ZDSYS201603311644527, the Longgang Key Laboratory Project under Grants No. ZSYS2017003 and No. LGKCZSYS2018000015, and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300701), Hefei National Laboratory, Hefei 230088, China. The authors would thank Ms. Floria Chen and Ms. Chuanyi Yang for their mentorship and revision tips. The authors would also thank Mr. Sixuan Mao for suggestions on frameworks of artificial neural networks.
@@
H.H. conceived the experiment and conducted the majority of the manuscript writing. H.H., Z.X. and Q.X. conducted the experiment and analyzed the results. Z.X. wrote the section on the principles of the neural network and prepared Figure 1. Q.X. wrote the section on the method of generation of dataset. Z.Z. provided guidance and funding support. All authors reviewed the manuscript.
Data availability statement
Data is provided within the manuscript or supplementary information files.